# Volatility Spillovers and Interconnectedness in OPEC Oil Markets: A Network-Based log-ARCH Approach

### v1

@@ Appendix A Forecast Sensitivity Analysis
Euclidean-Diss 3.9511 3.0325 4.4192 4.1742 3.9511 4.1391 3.9511 3.8108 3.8230
Correlation-Diss 4.1527 3.1918 4.6287 4.3773 4.1527 4.3557 4.1527 4.0089 4.0215
Piccolo-Diss 5.9245 5.3727 6.0643 6.0301 5.9245 6.0394 5.9245 5.8645 5.8698
Net CCC-GARCH 2.4196 1.8374 2.8355 2.6929 2.4196 2.5000 2.4196 2.3355 2.3428
Net DCC-GARCH 2.4196 1.8374 2.8356 2.6929 2.4196 2.4999 2.4196 2.3355 2.3427
Net GO-GARCH 2.4182 1.8373 2.8338 2.6930 2.4182 2.4996 2.4182 2.3347 2.3419
@@
Std. CCC-GARCH 2.6184 1.9018 2.6858 2.7018 2.6184 2.7358 2.6184 2.5723 2.5760
Std. DCC-GARCH 2.3854 1.6843 2.5488 2.5013 2.3854 2.4896 2.3854 2.3363 2.3403
Std. GO-GARCH 6.0758 5.6669 6.2848 6.2449 6.0758 6.1759 6.0758 6.0462 6.0484
Notes: Baseline uses $T_{0}=300$, $\varepsilon=\min Y^{2}_{\neq 0}$. Bold entries highlight the best performance in both Network-based models and Standard Multivariate GARCH models.
Appendix B Model Losses and MCS Membership
Figure 3: Model Losses and MCS Membership
Note: Models eliminated from the MCS include the network-based log-ARCH models with Euclidean-, Correlation-, and Piccolo-based dissimilarities, as well as the standard GO-GARCH model. Green bars indicate models retained in the 95% Model Confidence Set (MCS); gray bars were eliminated. The standard DCC-GARCH model achieved the lowest forecast loss and remained in the superior set. Network models based on CCC, DCC, and GO-GARCH weights remain statistically indistinguishable from the benchmark at the 5% significance level.

### v2

@@ Appendix E Forecast Sensitivity Analysis
Euclidean-Diss 2.7424 2.1112 3.2235 3.0711 2.8461 2.9360 2.7687 2.7424 2.7437
Correlation-Diss 2.7462 2.1166 3.1284 3.0369 2.7950 2.8476 2.7725 2.7462 2.7475
Piccolo-Diss 4.4773 3.9880 3.4166 3.5491 4.2584 6.9031 4.4984 4.4773 4.4783
Net CCC-GARCH 2.3267 1.8512 2.8274 2.6969 2.5425 2.5365 2.3550 2.3267 2.3281
Net DCC-GARCH 2.3268 1.8512 2.8259 2.6955 2.5426 2.5365 2.3552 2.3268 2.3282
Net GO-GARCH 2.3257 1.8514 2.8215 2.6932 2.5422 2.5349 2.3541 2.3257 2.3271
@@
Std. CCC-GARCH 2.5167 1.7831 2.7350 2.7583 2.8207 2.5669 2.5167 2.5167 2.5167
Std. DCC-GARCH 2.3796 1.6661 2.5726 2.5342 2.5439 2.4116 2.3796 2.3796 2.3796
Std. GO-GARCH 6.2708 5.9035 6.3449 6.3171 6.3152 6.1151 6.2656 6.2708 6.2626
Notes: Baseline uses $T_{0}=450$ and $\varepsilon$ set to the 1% quantile of non-zero squared returns. Bold entries indicate the best-performing model (lowest error) in each column. While Std. DCC-GARCH minimizes MAFE, the Net GO-GARCH model achieves the lowest RMSFE in the extended baseline window ($T_{0}=450$).
Appendix F Network Weights under Different In-Sample sizes
$T_{0}=200$
$T_{0}=250$
$T_{0}=300$
$T_{0}=350$
$T_{0}=400$
$T_{0}=450$
Figure F.1: Evolution of Network Weights for Different In-Sample Sizes ($T_{0}$). Note: Network plots of the six OPEC oil-exporting countries based on six different distance measures. Raw data-based distances: (a) Euclidean distance, (b) Correlation-based distance. Model-based distances: (c) Piccolo-based distance, (d) CCC-GARCH, (e) DCC-GARCH, and (f) GO-GARCH. The grey colour scale of the edges is proportional to the weight degree of the connections.